OpenStudy (mtalhahassan2):

Determine the derivative of each of the following:... f(x)= e^-x^3/x

OpenStudy (mtalhahassan2):

@Jamierox4ev3r

OpenStudy (mtalhahassan2):

so that be f1(x)= e^-x^3 . -x

zepdrix (zepdrix):

This is the function?\[\Large\rm f(x)= \frac{e^{-x^3}}{x}\]

OpenStudy (mtalhahassan2):

yes

zepdrix (zepdrix):

So you'll have to apply quotient rule, ya?

zepdrix (zepdrix):

Here is our quotient rule setup,\[\Large\rm f'(x)= \frac{\color{royalblue}{\left(e^{-x^3}\right)'}x-e^{-x^3}\color{royalblue}{\left(x\right)'}}{x^2}\]Derivatives need to be taken for the blue parts.

OpenStudy (mtalhahassan2):

f1(x)= -3x^2 e (x) - e^-x^3(1)/ (x)^2

OpenStudy (mtalhahassan2):

right??

jimthompson5910 (jim_thompson5910):

@MTALHAHASSAN2 you took the derivative of `e^(-x^3)` incorrectly

zepdrix (zepdrix):

Hmm, I'm just a little bit worried about your first e, did you forget to write the exponent?

OpenStudy (mtalhahassan2):

f1(x)= e^ -3x^2

OpenStudy (mtalhahassan2):

Is it be like that then

zepdrix (zepdrix):

Recall: \(\large\rm \dfrac{d}{dx}e^{x}=e^x\) You get the same exponential back. We need to simply apply the chain rule as an added step. \[\large\rm \frac{d}{dx}e^{-x^3}\quad=e^{-x^3}\frac{d}{dx}(-x^3)\quad=e^{-x^3}(-3x^2)\]

zepdrix (zepdrix):

You get the same exponential back (unchanged), plus some other stuff, due to the chain rule.

OpenStudy (mtalhahassan2):

oh ok

zepdrix (zepdrix):

times* some other stuff, due to the chain rule. that word plus is a lil misleading :) lol

zepdrix (zepdrix):

\[\Large\rm f'(x)= \frac{\color{royalblue}{\left(e^{-x^3}\right)'}x-e^{-x^3}\color{royalblue}{\left(x\right)'}}{x^2}\] \[\Large\rm f'(x)= \frac{\color{orangered}{\left(-3x^2e^{-x^3}\right)}x-e^{-x^3}\color{orangered}{\left(1\right)}}{x^2}\]

zepdrix (zepdrix):

And then simplify some stuff, ya? :)

OpenStudy (mtalhahassan2):

yeah

OpenStudy (mtalhahassan2):

are we simplify the x right

OpenStudy (mtalhahassan2):

f1(x)= -3x^3 e^-x^4

zepdrix (zepdrix):

Hmmm, I'm not sure what you did... Combine the x's in the first term,\[\Large\rm f'(x)= \frac{-3x^3e^{-x^3}-e^{-x^3}}{x^2}\]Then factor a -e^(-x^3) out of each term,\[\Large\rm f'(x)= \frac{-e^{-x^3}\left(~?~~+~~?~\right)}{x^2}\]What are you left with in the brackets? Should be two terms still.

OpenStudy (mtalhahassan2):

(-3x^2+1)

OpenStudy (mtalhahassan2):

right??

OpenStudy (mtalhahassan2):

@zepdrix

zepdrix (zepdrix):

Hmm, both terms were negative, so when you pull that negative out, they both become positive inside of the brackets, ya?

zepdrix (zepdrix):

(3x^2+1) Looks good besides that though :)

zepdrix (zepdrix):

Woops* (3x^3+1)

OpenStudy (mtalhahassan2):

but at the book of the book they have a different answer

OpenStudy (mtalhahassan2):